OpenStudy (anonymous):

y over negative 6 + 5 = 9. 24 -24 78 -78

ganeshie8 (ganeshie8):

\[\dfrac{y}{-6} + 5 = 9 \]

ganeshie8 (ganeshie8):

like this ?

OpenStudy (anonymous):

@ganeshie8 yes

ganeshie8 (ganeshie8):

good, its a two step equation our goal is to isolate \(y\)

ganeshie8 (ganeshie8):

start by subtracting `5` both sides

OpenStudy (anonymous):

@ganeshie8 ok

ganeshie8 (ganeshie8):

\(\dfrac{y}{-6} + 5 = 9\) subtract `5` both sides, you get : \(\dfrac{y}{-6} = 4\)

OpenStudy (anonymous):

yeah i got that @ganeshie8 and ok

ganeshie8 (ganeshie8):

next multiply `-6` both sides to isolate the \(y\) completely

OpenStudy (anonymous):

thank you soo much! @ganeshie8

ganeshie8 (ganeshie8):

\(\dfrac{y}{-6} + 5 = 9\) subtract `5` both sides, you get : \(\dfrac{y}{-6} = 4\) multiply `-6` both sides : \(y = -24\)

ganeshie8 (ganeshie8):

you should get `-24` as the final answer ^

OpenStudy (anonymous):

yeah! @ganeshie8